Solve the system by addition or substitution.
5x – 5y = 35
y = –7 + x

Answer by LilSkittleMd(119)

Solve the system by addition or substitution.
5x – 5y = 35
y = –7 + x
Substitution. Plug the y = equation into the y spot in the first equation.
5x - 5 (-7 + x) = 35
5x + 35 - 5x = 35
35=35
Since this is a true statement, the equations are the same line